A river 105.1 m wide flows toward the south at 33.6 m/min. A girl on the west bank wishes to reach the east bank in the least possible time. She can swim 105.1 m in still water in 1.70 min.

(c) What is her velocity relative to land?

Magnitude

direction south of east

(d) What is the total distance she travels?
(e) In what direction must she swim if she wishes to travel straight across the river?
Answer #1

flow speed = 33.6m/min = 0.56 m/s

swimming speed = 105.1m/1.7min = 1.03 m/s

(c) Net velocity = 1.03x - 0.56y

magnitude = 1.1727 m/s

direction atan(vy/ux) = 28.53 degrees south of east

(d) total time taken = 105.1/1.03 = 102.03 s

total distance = velocity amgnitude* time = 119.65 m

(e) direction as calculated above she must oppose the flow speed to swim straight

by swimming in directio 28.53 degrees north of east